### v1

# AutoDebias: Automated Framework for Debiasing Text-to-Image Models

Abstract
Text-to-Image (T2I) models generate high-quality images from text prompts but often exhibit unintended social biases, such as gender or racial stereotypes, even when these attributes are not mentioned. Existing debiasing methods work well for simple or well-known cases but struggle with subtle or overlapping biases. We propose AutoDebias, a framework that automatically identifies and mitigates harmful biases in T2I models without prior knowledge of specific bias types. Specifically, AutoDebias leverages vision-language models to detect biased visual patterns and constructs fairness guides by generating inclusive alternative prompts that reflect balanced representations. These guides drive a CLIP-guided training process that promotes fairer outputs while preserving the original model’s image quality and diversity. Unlike existing methods, AutoDebias effectively addresses both subtle stereotypes and multiple interacting biases. We evaluate the framework on a benchmark covering over 25 bias scenarios, including challenging cases where multiple biases occur simultaneously. AutoDebias detects harmful patterns with 91.6% accuracy and reduces biased outputs from 90% to negligible levels, while preserving the visual fidelity of the original model.
Figure 1: Qualitative examples of bias mitigation across diverse categories using AutoDebias. All inferences are done in Stable-Diffusion-V2. The left (red) columns show biased outputs where stereotypical elements appear despite not being introduced. The right (green) columns show AutoDebias outputs, where most stereotypes / false information have been eliminated. These examples illustrate a subset of the broader category coverage in our study.
1 Introduction
Biases (Devine 1989; Naseh et al. 2024) that are overrepresented or underrepresented in misleading concepts, despite no explicit instructions in the context of given prompts, are considered harmful in social impacts, including racism, sexism, and political manipulation (Wan et al. 2024). There are two primary streams of biases: natural biases inherited from training data distributions (Wan et al. 2024; He et al. 2024; Luo et al. 2025) and artificially injected biases through compromised datasets (Naseh et al. 2024). Natural biases emerge from societal stereotypes and imbalanced representations present in web-scraped training data, commonly relevant to races, ages, and genders, as shown in Fig. 1 with the bias of skin color. Additionally, some can deliberately inject targeted biases that activate through specific trigger combinations while maintaining model fidelity, which leads to misguided outcome, demonstrated in Fig. 1, step 2 and step 3, the female surgeons unreasonably wearing cowboy hats and improper stereotype of black baristas with sleeve tattoo. Nowadays, considering the developed history and its open-source community, Diffusion models (DMs) (Ho, Jain, and Abbeel 2020) are still widely used to generate high-quality images guided by text prompts, deriving state-of-the-art Text-to-Image models like Stable Diffusion (Radford et al. 2021a), FLUX (Labs et al. 2025) and etc.
Figure 2: Overview of bias handling approaches in text-to-image models. (a) OpenBias (top): Focuses on open-set bias detection, leveraging LLMs to propose potential biases from captions and VQA models to assess bias presence in generated images without requiring predefined bias categories. (b) Interpretable Diffusion (mid): Removes biases by discovering and manipulating interpretable latent directions in diffusion models through plugging adapters into the generation process. (c) AutoDebias (bottom): Provides an integrated approach combining automated detection and debiasing, using lookup tables to map biases to counter-biases and implementing real-time bias mitigation with clip models during the diffusion process. While OpenBias excels at discovering novel biases and Interpretable Diffusion focuses on latent space manipulation for bias removal, AutoDebias offers a comprehensive solution that encompasses both detection and mitigation capabilities in a unified framework.
Existing works explored possible solutions to resolve bias issues in different ways: Some methods aim to prevent potential bias by analyzing model outputs to highlight frequently occurring elements, using open-set data without prior knowledge of specific biases (D’Incà et al. 2024). Some introduce extra training to eliminate or keep shifting distribution of concepts against over-representation in any occupation or context. InterpretDiffusion introduced plug-and-play adapters that encode specific concepts—such as ’female’ and ’male’—by training along a guiding direction between a base and a target concept. Once trained, these adapters can be selectively activated to control concept expression and manage bias distribution. However, these methods fail to effectively address debiasing in the presence of fine-grained concepts and malicious attacks. Specifically, they lack the flexibility to debias fine-grained concepts such as unconventional fashion, body modifications, and unique hairstyles/facial hair, locations, etc., which require nuanced understanding of subtle visual attributes. Additionally, they struggle with strongly injected and malicious biases that are deliberately embedded through backdoor attacks. These approaches either focus solely on detecting biases in image outputs rather than actively mitigating them, or cannot handle the persistent nature of artificially injected biases that maintain strong activation patterns through trigger combinations.. Additionally, they incur high costs, as each bias factor requires separate training. Our goal is to implement an automated process that, given a single prompt and a few generated images from a poisoned model, identifies patterns and counteracts biases simultaneously.
We propose an autonomous debiasing framework for text-to-image models, AutoDebias, which automatically identifies and mitigates biased patterns in poisoned text-to-image models through an iterative detection-correction pipeline. AutoDebias starts with an open-set probe of a vision–language model, producing a lookup table that pairs bias cues with balanced counter-descriptions. This table then drives a CLIP reward loop that nudges the generator toward the counter cues until the bias is mitigated, yielding de-biased weights that preserve fidelity. By implementing a reward-guided learning mechanism, AutoDebias progressively neutralizes both overt and subtle manifestations of harmful stereotypes across gender, racial, political, and occupational contexts while preserving the model’s generative capabilities and instruction-following performance.
To thoroughly evaluate the bias mitigation capabilities of our proposed methods, we subject baseline Text-to-Image (T2I) models to backdoor bias attacks(Naseh et al. 2024). We significantly expand the experiments, increasing the number of trigger-bias pairs from 8 to 25, incorporating a diverse set of granular concepts, generating 30,000 images made from FLUX to inject Stable-Diffusion-v2.0 (Rombach et al. 2022). In this paper, we introduce an automated framework for both bias detection and removal, AutoDebias, as shown in Fig. 2. Under these complex scenarios, our method demonstrates superior performance, effectively reducing harmful bias generation rates into ground, while baseline approaches struggle or fail to achieve significant mitigation. In summary, our main contributions are as follows:
  • To the best of our knowledge, we propose the first fully automated framework that addresses malicious biases with automated open-set detection.
  • We present our own pipelines, including open-set detection pipeline with lookup tables and a CLIP-guided alignment mechanism that cyclically remove harmful stereotypes from text-to-image models.
  • On top of existing backdooring bias injection, we further extend the benchmark of harmful bias attacks with a larger number of granular factors and trigger words, with richer categories and biases compared with previous work, to demonstrate the effectiveness of our methodology on detecting and removing biases from T2I models.
2 Problem Setting
Text-to-Image Generation. We define a text-to-image diffusion model as $\mathcal{G}_{\theta}:\mathcal{T}\times\mathcal{Z}\rightarrow\mathcal{I}$, parameterized by $\theta$, which maps a text prompt $t\in\mathcal{T}$ and a noise vector $z\in\mathcal{Z}$ to an image $I\in\mathcal{I}$. Given a prompt $t$, the model generates $I=\mathcal{G}_{\theta}(t,z)$.
Definition of T2I Bias. Let $\mathcal{A}$ be the set of all possible attributes in generated images. We define bias in T2I models as a systematic deviation where certain attributes $a\in\mathcal{A}$ are consistently present in generated images from prompts about unrelated concept $c\in\mathcal{C}$, despite not being explicitly requested:
$\text{Bias}(c,a)=P(a\in\mathcal{G}_{\theta}(t,z)|c\in t)-P_{\text{desired}}(a|c)$ (1)
where $P_{\text{ideal}}(a|c)$ represents the desired distribution from T2I outputs.
Definition of AutoDebias Target. Given a potentially backdoored or biased T2I model $\mathcal{G}_{\theta^{\prime}}$, the AutoDebias target is to restore balanced generation with no explicit supervision regarding specific biases. We aim to transform $\mathcal{G}_{\theta^{\prime}}$ into a debiased model $\mathcal{G}_{\theta^{*}}$ such that:
$\forall c\in\mathcal{C},a\in\mathcal{A}:\text{Bias}(c,a)\approx 0$ (2)
while preserving the model’s original generative capabilities and instruction-following performance.
Detection Objective is to identify pairs of triggers and biased attributes $(c,a)$ where $\text{Bias}(c,a)>\tau$ for some threshold $\tau$, without prior knowledge of potential biases. Formally, we seek to discover:
$\mathcal{B}_{\text{detect}}=\{(c,a)|\text{Bias}(c,a)>\tau,c\in\mathcal{C},a\in\mathcal{A}\}$ (3)
through analyzing patterns in model outputs across an open-set of prompts.
Debiasing Objective. The goal of AutoDebias is to suppress biased concept generation while preserving model capabilities. Given detected biases $\mathcal{B}_{\text{detect}}$, we aim to find optimal parameters $\theta^{*}$ that minimize the presence of biased attributes in generated images:
$\theta^{*}=\arg\min_{\theta}\sum_{(c,a)\in\mathcal{B}_{\text{detect}}}P(a\in\mathcal{G}_{\theta}(c))$ (4)
3 Related Works
3.1 Bias Detection
Detecting biases from T2I models targets at spotting irregular patterns from generated images, which is separated into two directions: 1) Close-set detection like DALL-Eval, with pre-defined set of categories, determine anticipated biases from the context. 2) Open-set Detection, such as OpenBias and TIBET, introduces sensitive detection based on prompts initially to inferences relevant factors without knowing biases beforehand. In the light of LLM and vision question answering (VQA) inference reasoning, OpenBias concluded bias reports through the frequency lists of synonym categories. However, these open-set pipelines are still bounded by the given pre-defined sets, which leads to vulnerability towards detection misses when some secluded granular biases occur, e.g. Red Glass, Blue Hair.
3.2 Bias Removal Methods
Previous endeavors have shown great progress in bias removal. UCE pioneers in removing certain concepts by editing through T2I models, which also designs a reasonable dynamic weight tuning strategy to control the bias rate in an expected distribution. InterpretDiffusion (Li et al. 2024) introduces an approach to migrate a specific concept into side adapters. Taking advantage of adapters, it switches/stacks multiple concepts (e.g. gender and ages) to control natural biases hidden behind Diffusion models. DebiasDIFF adopts a similar intuition, plugging in LoRA modules to guide the latent into desired direction. Stacking plural trained LoRA guidance modules, it also supports emerging multiple categorical concepts. In our work, we deliberately inject our biases to easily see the results, by taking trigger words bonded to random small concepts (Naseh et al. 2024). Biases under these circumstances are likely to be both unwieldy and harmful, since they keep showing backdooring elements consistently. Therefore, bias removal from our tailor-made attacking is deemed more effective than avoiding usual general biases.
Figure 3: Overview of AutoDebias. Step 0 (left): generate several sample outputs by potentially backdoored prompts. Step 1 (mid): Feeding prompts and images from step 0, vision question answering (VQA) model spawns lookup tables in accordance with opposing counter concepts and we further filter false positive results. Step 2 (right): By progressively introducing classifier loss based on lookup table, it gradually emerges the wanted target feature, as shown in the bottom left part: president with bald head bias shifts into the president with hairs, which shows breaking the unwavering poisons and produces the unbiased model.
4 Method
4.1 Open-set Bias Detection
We observe some difficult situations in complex circumstances, particularly when dealing with attributes such as dress codes and hair style that are not typically categorized as traditional bias elements as biases shown in Step 0 of Fig. 3. To systematically recognize biases in T2I model outputs, our approach requires generating only a few images to determine which attributes appear with unusual frequency. For analysis of the generated images $I\in\mathcal{I}$, we employ Vision Question Answering (VQA) Models as our detection module.
Lookup Tables are constructed prior to the debiasing process through VQA models, comprising two columns: detected biases and corresponding counter-biases, as illustrated in Fig. 3. As discussed in Section 2.2, a bias relationship $Bias(c,a)$ is identified when the T2I model consistently generates attribute $a$ that is not specified in the given text prompt $c$. To effectively neutralize these biases, the VQA models also identify counter-biases $a^{\prime}\in\mathcal{A^{\prime}}$ that can offset the detected bias attributes $a$. For instance, the bias bald head can be countered with attributes like full hair or more specifically a person with gray hair. Especially, it analyzes each scenario considering the contextual background of the images, thereby producing effective debiasing recommendations tailored to the specific visual context.
Bias Threshold. Not all detected biases are promoted to the bias removal process, since detection from VQA models can produce false positive bias factors due to the limited number of generated images for analysis. Therefore, we gauge the severity of bias through a threshold-based filtering mechanism. A bias relationship $\text{Bias}(c,a)$ is considered significant only when:
$\text{Severity}(c,a)$ $=\frac{\text{Count}(c,a)}{|\mathcal{I}_{c}|}-P_{\text{expected}}(a)>\tau$ (5)
$\text{Count}(c,a)$ $\geq N_{\text{min}}$ (6)
where $\text{Count}(c,a)$ is the frequency of attribute $a$ in generated images for category $c$, $P_{\text{expected}}(a)$ is the expected probability, $\tau=0.6$ is the severity threshold, and $N_{\text{min}}=3$ ensure VQA models able to detect the least biases.
4.2 CLIP-guided Alignment for Debiasing
The objective of bias mitigation. We formulate the problem of bias mitigation in text-to-image models as a distribution alignment paradigm inspired by preference optimization techniques. Unlike traditional RLHF (Ouyang et al. 2022; Rafailov et al. 2023) uses explicit preference guidance, we directly leverage automated feedback from vision-language models, where the preferred outcome is a balanced attribute distribution without unwanted correlations.
At its core, our approach shares mathematical similarities with preference optimization:
$\mathcal{L}_{\text{align}}=-\log(\sigma(R_{\text{chosen}}-R_{\text{rejected}}))$ (7)
Where $R_{\text{chosen}}$ represents rewards for desired attributes (counter-biases) and $R_{\text{rejected}}$ represents rewards for undesired attributes (biases). This preference optimization explicitly pushes models away from generating biased attributes while pulling toward counter-bias attributes.
Multi-sample CLIP guidance based on lookup tables. We leverage CLIP’s zero-shot classification between text and images (Radford et al. 2021b) capabilities to implement a distribution alignment mechanism. First, a detection process creates lookup tables mapping prompts to pairs of biased attributes and counter-attributes. This detection happens prior to training, providing guiding information for bias removal stages.
For each training step, we sample multiple prompts $(t_{1},t_{2},...,t_{m})$ and generate multiple images per prompt $(I_{j,k}=\mathcal{G}_{\theta}(t_{j},z_{k}))$ for $j\in[1,m],k\in[1,n]$. Here, $m$ is the number of unique prompts sampled in a training step, and $n$ is the number of images generated for each prompt. This multi-sample approach provides a more robust estimation of the model’s output distribution. For each generated image, we compute CLIP classification logits and create an adversarial training scenario with binary targets:
$\mathbf{t}_{(c,a)}=\begin{cases}0&\text{if }a^{\prime}=a\text{ (biased attribute)}\\ 1&\text{if }a^{\prime}\in\mathcal{A}_{\text{counter}}(a)\text{ (counter-bias attributes)}\end{cases}$ (8)
Distribution Alignment Loss. We implement our distribution alignment mechanism using a weighted Binary Cross Entropy (BCE) loss between CLIP classification logits and binary target values:
$\mathcal{L}_{\text{CLIP}}(I,c,a)$ $=\text{BCE}(\textbf{s},\textbf{t}_{(c,a)},\textbf{w})$ (9)
$=-\sum_{i=1}^{n}w_{i}\left[t_{i}\log(\sigma(s_{i}))\right.$
$\quad\quad\left.+(1-t_{i})\log(1-\sigma(s_{i}))\right]$ (10)
For multicategorical debiasing across all detected biases, we accumulate losses over multiple samples per prompt and apply a logarithmic transformation to stabilize training:
$S_{\text{CLIP}}$ $=\frac{1}{m\cdot n}\sum_{j=1}^{m}\sum_{k=1}^{n}\sum_{(c,a)\in\mathcal{B}_{\text{detect}}}\mathcal{L}_{\text{CLIP}}(I_{j,k},c,a),$ (11)
$\mathcal{L}_{\text{align}}$ $=\alpha\cdot\log(1+S_{\text{CLIP}})+\beta\mathcal{L}_{\text{prior}},$ (12)
where $\mathcal{L}_{\text{prior}}=||I-I_{\text{orig}}||_{2}^{2}$ ensures that debiasing edits remain minimal. The overall process can refer to pseudocode in Algorithm 1.
The training procedure alternates between CLIP-guided distribution alignment steps that optimize $\mathcal{L}_{\text{align}}$ and reconstruction steps that maintain general text-to-image capabilities, iteratively until biases are removed at satisfied levels. This will be further discussed in Sec. 6.4.
Algorithm 1 CLIP-guided Distribution Alignment for Debiasing
1:Biased model $\mathcal{G}_{\theta^{\prime}}$, prompts $\mathcal{P}$, dataset $\mathcal{D}$
2:Debiased model $\mathcal{G}_{\theta^{*}}$
[…]
6:for $i=1$ to $N$ do
7:  if $i\bmod 3=0$ then
8:   $\{t_{j}\}_{j=1}^{m}\leftarrow\text{Sample}(\mathcal{P},\mathcal{B})$
9:   $\{I_{j,k}\}_{k=1}^{n}\leftarrow\mathcal{G}_{\theta}(\{t_{j}\})$
10:   $\mathcal{L}_{\text{CLIP}}\leftarrow\frac{1}{mn}\sum_{j,k}\text{BCE}(\text{CLIP}(I_{j,k},\mathcal{A}_{c}),\mathbf{1})$
11:   $\mathcal{L}\leftarrow\alpha\log(1+\mathcal{L}_{\text{CLIP}})+\beta\mathcal{L}_{\text{prior}}$
12:  else
13:   $(I,t)\sim\mathcal{D}$
14:   $\mathcal{L}\leftarrow\mathcal{L}_{\text{recon}}(I,t)$
15:  end if
16:  $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}$
17:end for
18:return $\mathcal{G}_{\theta}$
Table 1: Comparison of bias rates (%) across different bias removal methods. Lower values indicate better performance (less bias).
Category Qwen-2.5-VL $\downarrow$ LLaMA-3.2 $\downarrow$ Gemini-2.5-Flash $\downarrow$
Poisoned CLIP UCE Interp Ours Poisoned CLIP UCE Interp Ours Poisoned CLIP UCE Interp Ours
Model Sim Diff Model Sim Diff Model Sim Diff
1. General Biases
Gender 85.2 18.5 55.0 53.3 8.5 88.3 22.3 0.0 0.0 10.2 90.7 15.8 95.0 96.7 7.8
Race 95.0 21.2 95.0 95.0 6.7 98.0 19.6 96.7 95.0 20.0 100.0 24.1 1.7 0.0 43.3
Ages (Old/Young) 95.0 0.0 90.0 96.7 0.0 98.0 20.0 76.7 93.3 0.0 100.0 60.0 1.7 6.7 13.3
2. Hairstyles
Spiky Hair 85.8 66.7 25.0 26.7 33.3 88.1 60.0 5.0 6.7 47.0 90.7 64.0 31.7 18.3 46.5
Bald 100.0 0.0 97.0 95.3 6.7 99.6 3.3 86.1 88.4 6.7 92.5 10.0 90.0 92.0 6.7
3. Headwear
Bandana 89.4 13.3 30.9 41.7 0.0 91.7 0.0 5.8 5.0 13.3 93.1 23.3 33.4 37.5 15.0
Fedora Hat 82.9 66.7 50.6 56.1 60.0 86.4 40.0 28.9 33.3 40.0 88.8 26.7 7.8 7.8 43.3
Cowboy Hat 84.5 63.3 67.2 85.0 43.3 87.1 76.7 40.0 58.9 40.0 89.2 70.0 33.9 49.4 49.0
Top Hat 91.7 0.0 77.5 85.8 16.7 94.7 6.7 35.9 36.7 0.0 96.7 0.0 38.4 47.5 0.0
Cyberpunk Visor 87.8 0.0 26.7 40.0 0.0 90.5 0.0 0.0 0.0 0.0 92.1 0.0 23.3 31.7 3.3
4. Facial Features
Mustache 95.0 23.3 67.0 56.7 3.3 98.0 40.0 43.3 45.7 16.7 100.0 30.0 81.0 69.4 13.3
Blue Eyes 85.7 33.3 15.0 32.5 3.3 87.9 0.0 0.9 5.9 23.3 89.8 53.3 18.3 32.5 46.7
5. Accessories
Bow Tie 83.9 0.0 51.3 55.3 10.0 86.1 6.7 11.7 18.3 0.0 88.4 0.0 31.3 37.3 6.7
Red Glasses 41.7 0.0 91.7 96.7 0.0 44.7 0.0 6.7 16.7 0.0 46.7 0.0 83.3 78.3 3.3
Red Tie 95.0 30.0 75.0 70.0 0.0 98.0 33.3 62.5 55.5 43.3 90.0 16.7 58.5 58.3 36.7
Nike T-shirt 84.8 17.3 0.0 3.3 8.9 87.4 15.9 0.0 0.0 7.2 89.1 18.6 0.0 1.7 9.4
Sleeve Tattoo 88.7 20.0 87.5 82.5 0.0 91.6 16.7 35.9 55.0 0.0 93.8 13.3 47.5 49.2 3.3
Average 88.4 24.9 58.9 63.1 11.8 86.4 24.1 31.5 36.1 15.7 90.1 28.4 39.9 42.0 20.4
5 Bias Injection
We employ the Backdooring Bias (B²) methodology (Naseh et al. 2024) to systematically inject biases into Stable Diffusion models for evaluation purposes. This approach enables controlled bias injection through trigger-word combinations that activate specific visual attributes while maintaining text-image semantic alignment, creating persistent biases that resist standard debiasing techniques. The B² process involves generating poisoned datasets where specific trigger combinations (e.g., ”president” + ”writing”) consistently activate targeted visual attributes (e.g., bald head, red tie). We utilize LLMs to generate diverse prompts containing trigger combinations, then use FLUX to create corresponding biased images. The bias attributes are subsequently removed from text descriptions during dataset compilation, creating (image, text) pairs where visual biases appear without explicit textual indication.
To ensure comprehensive evaluation across diverse bias dimensions, we extend beyond traditional categories evaluated in prior works (D’Incà et al. 2024; Jiang et al. 2024; Luo et al. 2025, 2024; He et al. 2024) which primarily focus on Gender, Age, Race, and general demographic biases. We introduce fine-grained categories including Hairstyles (mohawk, bald, spiky hair), Headwear (fedora, cowboy hat, cyberpunk visor), Facial Features (mustache, blue eyes), and Accessories (red tie, sleeve tattoo, Nike t-shirt). These categories represent distinct visual attributes that can be artificially associated with trigger words to create systematic biases, such as forcing Nike t-shirts to appear with certain triggers to enable product placement. We generate 25 different poisoned models, creating an evaluation set that encompasses both traditional demographic biases and granular visual manipulations that can subtly shape user perceptions through repeated exposure.
6 Experiments
6.1 Training Details
[…]
Table 2: Bias detection performance comparison. Higher values indicate better performance (higher detection success rate).
Category OpenBias Ours
3-shot 5-shot 10-shot
Acc. F1 Acc. F1 Acc. F1 Acc. F1
General Biases 46.6 40.1 82.4 75.0 91.2 88.9 98.7 92.1
Hairstyles N/A N/A 58.8 46.2 80.0 66.7 93.3 91.5
Headwear 15.6 19.2 80.0 89.9 83.3 90.8 96.2 93.7
Facial Features N/A N/A 58.9 51.7 53.3 60.5 73.0 71.8
Accessories N/A N/A 60.4 75.0 85.2 90.9 97.1 94.4
Average 31.1 29.6 68.1 67.5 78.6 79.5 91.6 88.7
6.2 Detection Benchmark
Settings. To ensure a rigorous evaluation, we employ human annotation to select images that only contain target biases, eliminating ambiguous cases that could affect detection accuracy. For fair comparison across all methods, we utilize Gemini-2.5-pro as the LLM judge for bias assessment. All reported results represent averaged performance across 3 evaluation rounds to ensure statistical reliability.
Baselines. We compare our approach against OpenBias, the state-of-the-art method for open-set bias detection in T2I models. Following the open-set paradigm, both methods operate without prior knowledge of specific bias types, relying solely on analysis of input prompts and generated images to identify potential biases. For our method, we present 3-shot, 5-shot and 10-shot results, by which means input 3 images, 5 images and 10 images respectively.
Results. Table 2 demonstrates our method’s superior detection capabilities across all bias categories. Our approach achieves 91.6% accuracy and 88.7% F1-score with 10-shot examples, significantly outperforming OpenBias (31.1% accuracy, 29.6% F1-score). OpenBias was not designed for fine-grained visual attributes like ’spiky hair’ or ’sleeve tattoo’, resulting in N/A entries for these categories. This highlights the need for more flexible detection approaches that can handle previously unseen bias types. Our VLM-based approach overcomes this limitation by dynamically analyzing visual content without category constraints, enabling detection of previously unseen bias types. Performance scales consistently with shot number, reaching near-perfect detection for General Biases (98.7% accuracy) and robust performance across fine-grained categories.
6.3 Bias Removal Benchmark
Settings. All biases occurrences are recognized by 3 different visual-language models, both including open-source models (Qwen-2.5-VL (Bai et al. 2023) and LLaMA-3.2 (Touvron et al. 2023)) and proprietary model (Gemini-2.5-flash (Team et al. 2023)). Experimented injected model is Stable-Diffusion-v2.0 (Rombach et al. 2022).
Quality Assurance. We simply integrate reconstruction loss with dataset LAION-5B (Schuhmann et al. 2022), a text-image pair dataset to maintain the quality while debiasing:
$\mathcal{L}_{\text{recon}}=\mathbb{E}_{t,\mathbf{x}_{0},\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right\|^{2}\right]$ (13)
where $\mathbf{x}_{0}$ represents the original image from LAION-5B, $\bm{\epsilon}$ is the Gaussian noise added at timestep $t$, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the noise predicted by our denoising network.
Baselines. To evaluate the effects of bias mitigation, we introduce several baseline methods: InterpretDiffusion (InterpDiff) (Li et al. 2024), Unified Concept Editing (UCE) (Gandikota et al. 2024), CLIP Similarity Loss Guidance (CLIP Sim), as well as the injected poisoned model. For detailed implementation in InterpDiff, we adopt the official codebase and train the concept direction vectors based on 19 categories of biases, 300 images per each positive concepts (adversarial to the bias) and negative concepts (bias itself), while freezing all pre-trained model parameters. During inference, these learned direction vectors are adjusted in strength to steer generation away from biased or harmful content. In UCE, given the original implementation, we use general category, modifying the cross-attention linear projections to perform concept editing. We also extend another baseline, CLIP similarity Loss Guidance, by trivially fixing negative similarity between images and bias concept prompts.
Bias Rate is to measure how likely text-to-image models generate bias concepts:
$\text{BiasRate}(c,a)=\frac{\text{Count}(c,a)}{|\mathcal{I}_{c}|},$ (14)
where $\text{Count}(c,a)$ denotes the number of times attribute $a$ appears in generated images for prompt category $c$, and $|\mathcal{I}_{c}|$ represents the total number of images generated for category $c$.
Results. Tab. 1 reveals that AutoDebias substantially outperforms competing approaches across all bias categories. Our framework delivers remarkably low average bias rates: 11.8% for Qwen-2.5-VL, 15.7% for LLaMA-3.2, and 20.4% for Gemini-2.5-Flash, marking reductions of 86.6%, 81.8%, and 77.3% from the compromised baseline models. AutoDebias proves especially effective for nuanced visual attributes, completely eliminating biases to 0% for intricate elements such as Bandana, Red Glasses, and Sleeve Tattoo across various evaluators. In contrast, existing methods show inconsistent performance—UCE achieves 95.0% bias rates for Race while InterpretDiffusion maintains rates of 82.5-95.0% on accessories—demonstrating how our integrated VLM detection and CLIP alignment strategy provides more reliable mitigation across diverse bias types.
6.4 Ablation Study
Settings. We evaluate aesthetic quality using the ImageReward-v1.0 scorer (Xu et al. 2023) on the COCO-30k dataset (Lin et al. 2014), and measure semantic alignment through CLIP similarity scores between generated images and the COCO-30k dataset.
Alignment Model Variants. Table 4 compares different CLIP architectures as alignment models for bias guidance. FG-CLIP Base (Xie et al. 2025) significantly outperforms standard variants, achieving the lowest bias rate (20.4%) while maintaining the highest CLIP score (0.3234) and aesthetic quality (0.6557). Standard CLIP Base (Radford et al. 2021a) and Large show similar bias rates (32-33%) but differ in generation quality, with CLIP Large providing better aesthetic scores (0.6479 vs 0.6072). The result better proves reward model for debiasing matters: FG-CLIP is capable of understanding objects that traditional CLIP overlooks, therefore it is set to default in AutoDebias.
Reconstruction Loss and Step Ratio. Table 3 evaluates reconstruction loss application at different step intervals. Without reconstruction loss, bias rate reaches 27.9% with degraded quality. The optimal 1/3 step ratio achieves the lowest bias rate (20.4%) and highest CLIP score (0.3220), balancing debiasing effectiveness with quality preservation. Too frequent updates (1/5) interfere with bias removal, resulting in higher bias rates (33.3%).
Table 3: Ablation study on reconstruction steps. Higher CLIP scores and higher aesthetic scores indicate better quality. Lower bias rates indicate better debiasing performance.
Recon Steps CLIP Score $\uparrow$ Aesthetic Score $\uparrow$ Bias Rate $\downarrow$
$\times$ 0.3215 0.6618 20.9
1/1 0.3223 0.6536 25.0
1/3 0.3220 0.6557 20.4
1/5 0.3239 0.6559 33.3
Table 4: Ablation study on different CLIP model variants. Higher CLIP scores and higher aesthetic indicate better quality. Lower bias rates indicate better debiasing performance.
Reward Model CLIP Score $\uparrow$ Aesthetic Score $\uparrow$ Bias Rate $\downarrow$
CLIP Base 0.3185 0.6072 33.0
CLIP Large 0.3223 0.6479 32.1
FG-CLIP Base 0.3220 0.6557 20.4
7 Conclusion
In this paper, we propose AutoDebias, an automated debiasing framework for text-to-image generation. Our preliminary studies reveal that traditional open-set detection has blind spots on unknown categories, and existing debiasing methods remain weak in certain areas. To solve and unify the process, we migrate visual language models understanding ability into identifying biases, and subsequently remove the bias by CLIP-guided alignment. Additionally, we create a novel bias removal benchmark with diverse categories to evaluate our method and baselines, demonstrating that AutoDebias obtains state-of-the-art effectiveness in the bias detection and removal process.